OpenStudy (anonymous):

Find the equation of the line that bisects the obtuse angles between the lines 11x+2y-7=0 and x+2y+2=0.

Parth (parthkohli):

There are two kinds of bisectors. One bisecting the acute angle, and the other bisecting the obtuse one.\[\dfrac{a_1 x + b_1 y + c_1}{\sqrt{a_1 ^2 + b_1^2}} = \pm \dfrac{a_2 x + b_2 y + c_2}{\sqrt{a_2 ^2 + b_2 ^2}}\]The above will give you the bisectors for \(a_1 x + b_1 y + c_1 = 0\) and \(a_2 x + b_2 y + c_2 = 0\). The underlying reason is that for any point \((x,y ) \) on the bisector, the distance of the point from both lines would be equal. In the next post, I'll explain how you can distinguish between the two bisectors...

Parth (parthkohli):

To distinguish between the two bisectors, follow these steps: 1. Make \(c_1\) and \(c_2\) positive. Case 1: If \(a_1 a_2 + b_1 b_2 > 0\), then the "+" sign will give you the obtuse angle bisector. Case 2: If \(a_1 a_2 + b_1 b_2 < 0\), then the "-" sign will give you the obtuse angle bisector.

OpenStudy (anonymous):

So what values should I plug in? ...

Parth (parthkohli):

1. Make \(c_1\) and \(c_2\) positive for both the lines.

OpenStudy (anonymous):

So that would be 11x+2y+7=0?

Parth (parthkohli):

Somewhat.\[11x + 2y - 7 = 0\]You're allowed to manipulate this in such a way that you make the constant term positive. I'm sure you know how you can do that.

OpenStudy (anonymous):

-11x-2y+7=0?

Parth (parthkohli):

Correct! Step 2: Check the sign of \(a_1 a_2 + b_1 b_2\). Careful with the values though. \(a_1 \) is not the coefficient of \(x\) in the original equation, but the new one.

OpenStudy (anonymous):

It's negative..

Parth (parthkohli):

Great, so the "-" sign will give you the equation of your bisector.\[\dfrac{\color{blue}{-11}x \color{blue}{-2}x + 7}{\sqrt{(-11)^2 + (-2)^2}} = - \dfrac{x + 2y + 2}{\sqrt{(1)^2 + (2)^2}}\]

OpenStudy (anonymous):

x=5y/4+3/2

Parth (parthkohli):

I'm not gonna check that, but I'm just gonna assume you're correct. =_= Good going!

OpenStudy (anonymous):

LMAO thank you! Can I ask you a few more questions?

Parth (parthkohli):

Sure!

OpenStudy (anonymous):

Find the equation of the line passing through (6,9) and is parallel to the line whose inclination is arctan (2).

Parth (parthkohli):

Hint: If the inclination is \(\theta\), then the slope is \(\tan \theta\).

Parth (parthkohli):

Oh, BTW, in the previous question, did you understand how that algorithm came into existence?

OpenStudy (anonymous):

Encountered that for the first time to be honest.. hehe

Parth (parthkohli):

OK, I'm gonna give you a full derivation of that once we're done with this question if you have the time for it. I was too lazy to type all of it out the last time.

OpenStudy (anonymous):

If I already got the slope I can now use point slope to get the equation right??

Parth (parthkohli):

Exactly.

OpenStudy (anonymous):

Okay thank you! :)

Parth (parthkohli):

Before I start with the derivation, how'd you have done the question? What method was taught to you?

OpenStudy (anonymous):

We have to graph the line..

Parth (parthkohli):

Yeah, but you need data to go with it too. Graphing the line would simply give you the way it'd look. I think you can manage it with geometry too, but anyway, since we've begun already...

Parth (parthkohli):



Parth (parthkohli):